OpenStudy (anonymous):

The spinning teacup ride at the fair makes 19 rotations in 2 minutes. How many rotations does it make in 5 minutes? 9.5 22 38 37.5

OpenStudy (anonymous):

\[\frac{2}{19} = \frac{5}{x}\]

OpenStudy (mathteacher1729):

How many revolutions does it make in one minute?

OpenStudy (anonymous):

When you cross multiply, it becomes \[2x = (19)(5)\]Can you figure out the rest?

OpenStudy (anonymous):

YES THAX

OpenStudy (anonymous):

np :)